# FusionDTI: Fine-grained Binding Discovery with Token-level Fusion for Drug-Target Interaction

### v3

@@ 4.3 Comparison of the BAN and CAN Fusion Modules
4.4 Efficiency Analysis
Efficiency in computational models is crucial, particularly when handling large-scale and extensive datasets in drug discovery. Our proposed model stores drug representations and target representations in memory for later online training. As evidenced by Figure 4, FusionDTI-CAN and FusionDTI-BAN with pre-encoded representations process the BindingDB dataset much faster than the non-pre-coded models, approximately 45 minutes and 220 minutes, respectively. This stark difference highlights the advantage of pre-encoded, which eliminates the need for real-time data processing and accelerates the overall throughput. While FusionDTI-BAN and DrugBAN have the same fusion module, the pre-encoded FusionDTI-BAN runs faster and predicts more accurately, as shown in Table 1. In addition, FusionDTI-BAN runs faster than FusionDTI-CAN, indicating that the BAN fusion module is more efficient. Ultimately, FusionDTI-BAN with pre-encoded data stands out as a highly efficient approach, offering substantial benefits in scenarios where exists large-scale data. We further analyse the time complexity in Appendix 6.6.
@@ 4.5 Ablation Study
| Aggregation | AUC | AUPRC | Accuracy |
| CLS | 0.982 | 0.983 | 0.956 |
| Pooling | 0.989 | 0.990 | 0.961 |
The fine-grained interaction of drug and target representations is critical in DTI as it directly impacts the model’s ability to infer potential binding sites. For FusionDTI, this interaction is facilitated by the CAN module, which markedly enhances the predictive accuracy by capturing the fine-grained interaction information between the drugs and targets. Table 4 demonstrates the impact of the CAN module on the prediction performance using the BindingDB dataset. When the fusion module is omitted, the model achieves an AUC of 0.954 and an accuracy of 0.894. Conversely, using the CAN module, there is a significant improvement, with the AUC increasing to 0.989 and the accuracy reaching 0.961. This highlights the effectiveness of the CAN module in improving the inference ability of FusionDTI. Additionally, in Table 4, we compare the performance of two aggregation strategies within the CAN module. The pooling strategy outperforms the CLS-based aggregation, achieving an AUC and AUPRC of 0.989 and 0.990, respectively. This comparison highlights the superior effectiveness of the pooling in aggregating contextual information. Thus, the integration of a CAN module, particularly employing a pooling aggregation strategy, is shown to be essential for making confident and accurate predictions.
@@ 4.6 Analysis of Fusion Scales
4.7 Evaluation of PLMs Encoding
The protein encoder and drug encoder are fundamental for the token-level fusion of representations, as these encoders are responsible for generating fine-grained representations to better explore interaction information. Our proposed model employs two PLMs encoding two biomedical entities: the drug and protein, respectively. In terms of the protein encoders, Figure 7 compares the the performance of the two protein encoders (SaProt [34] and ESM-2 [23]) in combination with three different drug encoders: ChemBERTa-2 [1], SELFormer [47] and MoLFormer [32]. From the figure, we find that SaProt consistently outperforms ESM-2 when combined with all three drug encoders. As can be seen in Figure 7, SELFormer achieves the best performance in encoding the drug sequences among the three advanced drug encoders. Notably, the top-performing combination is SaProt and SELFormer, hence our proposed FusionDTI uses them as drug and protein encoders.
@@ 4.8 Case Study
| 3. water molecules - Tyr129, Met132, Trp274, Try308 and Lys214; |
A further strength of FusionDTI to enable explainability, which is critical for drug design efforts, is the visualisation of each token’s contribution to the final prediction through cross-attention maps. To compare with the DrugBAN model, we examine three identical pairs of DTI from the Protein Data Bank (PDB) [4]: (EZL - 6QL2 [18], 9YA - 5W8L [29] and EJ4 - 4N6H [13]). As shown in Table 5, our proposed model predicts additional binding sites (in bold) evidenced by PDB [4] in comparison to the DrugBAN model. For instance, to predict the interaction of the drug EZL with the target 6QL2, our proposed model using BertViz [39] highlights potential binding sites as illustrated in Figure 8. Specifically, our CAN module is effective in capturing fine-grained binding information at the token level, as we have successfully predicted the novel binding between Gln92 and the benzothiazole ring [11]. In particular, we address the lack of structural information on protein sequences by employing the SA vocabulary, which matches each residue to a corresponding 3D feature via Foldseek [36]. This study highlights the effectiveness of FusionDTI in enhancing performance on the DTI task, thereby supporting more targeted and efficient drug development efforts. In Appendix 6.5, we present three pairs of prediction visualisations.
@@ 5 Conclusions
With the rapid increase of new diseases and the urgent need for innovative drugs, it is critical to capture and gauge fine-grained interactions, since the binding of specific drug atoms to the main amino acids is key to the DTI task. Despite some achievements, fine-grained interaction information is not effectively captured. To address this challenge, we introduce FusionDTI uses token-level fusion to effectively obtain fine-grained interaction information between drugs and targets. Through experiments on three well-known benchmark datasets, we demonstrate that our proposed FusionDTI model outperforms eight state-of-the-art baselines in DTI prediction, particularly in the more realistic cross-domain scenario. Additionally, we show that the attention weights of the token-level fusion module in our model can highlight potential binding sites, providing a certain level of explainability.
Limitations: Even if our proposed model identifies potentially useful DTI, these predictions need to be validated by wet experiments, a time-consuming and expensive process.

### v4

@@ 4.4 Ablation Study
The fine-grained interaction of drug and target representations is critical in DTI as it directly impacts the model’s ability to infer potential binding sites. For FusionDTI, this interaction is facilitated by the CAN module, which markedly enhances the predictive accuracy by capturing the fine-grained interaction information between the drugs and targets. Table 3 demonstrates the impact of the CAN module on the prediction performance. When the fusion module is omitted, the model achieves an AUC of 0.954 and an accuracy of 0.894. Conversely, using the CAN module, there is a significant improvement, with the AUC increasing to 0.989 and the accuracy reaching 0.961. This highlights the effectiveness of the CAN module in improving the inference ability of FusionDTI. In Appendix A.7 and A.8, we further compare time-consuming and time complexity with baselines.
@@ 4.6 Case Study
| 3. water molecules - Tyr129, Met132, Trp274, Tyr308 and Lys214; |
A further strength of FusionDTI to enable explainability, which is critical for drug design efforts, is the visualisation of each token’s contribution to the final prediction through cross-attention maps. To compare with the DrugBAN model, we examine three identical pairs of DTI from the Protein Data Bank (PDB) (Berman et al., 2007): (EZL - 6QL2 (Kazokaitė et al., 2019), 9YA - 5W8L (Rai et al., 2017) and EJ4 - 4N6H (Fenalti et al., 2014)), which are excluded from the training data. As shown in Table 4, our proposed model predicts more binding sites existing in the PDB (Berman et al., 2007) (in bold) by ranking the binding sites shown in the attention map. For instance, to predict the interaction of the drug EZL with the target 6QL2, our proposed model using BertViz (Vig, 2019) highlights potential binding sites as illustrated in Figure 5. Specifically, our CAN module is effective in capturing fine-grained binding information at the token level, as we have successfully predicted the novel binding between Gln92 and the benzothiazole ring (Di Fiore et al., 2008). In particular, we address the lack of structural information on protein sequences by employing the SA vocabulary, which matches each residue to a corresponding 3D feature via Foldseek (Van Kempen et al., 2024). This study highlights the effectiveness of FusionDTI in enhancing performance on the DTI task, thereby supporting more targeted and efficient drug development efforts. In Appendix A.9, we further investigate ten DTI pairs in non-small cell lung cancer (NSCLC) from PDB (Waliany et al., 2025), highlighting predicted binding residues.
@@ 5 Conclusions
With the rapid increase of new diseases and the urgent need for innovative drugs, it is critical to capture fine-grained interactions, since the binding of specific drug atoms to the main amino acids is key to the DTI task. Despite some achievements, fine-grained interaction information is not effectively captured. To address this challenge, we introduce FusionDTI uses token-level fusion to effectively obtain fine-grained interaction information. Through experiments on three well-known datasets, we demonstrate that our proposed FusionDTI model outperforms eight state-of-the-art baselines, particularly in the more realistic cross-domain scenario. Additionally, we show that the attention weights of the token-level fusion module can highlight potential binding sites, providing a certain level of explainability.
Limitations